OpenStudy (anonymous):

d 7th n 9th term of an ap are 346 n 328 find d 2nd term

OpenStudy (anonymous):

Just to restate your question. In the arithmetic progression 7th=346 and 9th=328. Need to find the 2nd term.

OpenStudy (hoblos):

9th= 7th +2d 328=346 +2d then d = -9 7th = 2nd + 5d 2nd = 7th - 5d = 346 + 45 = 391

OpenStudy (anonymous):

a7=a+(7-1)d d=-9 a7=346 346=a+6*(-9)=a-54 a=346+54=400 a2=400-9=391